OpenStudy (anonymous):

Mr. Edge needs 5 students, who will be arranged from tallest to shortest, for a demonstration. He has 21 students in his class. How many different combinations of students are possible for the demonstration?

OpenStudy (cwrw238):

21C5 = 21! ----- 16! 5!

OpenStudy (cwrw238):

or you can write it as 21*20*19*18*17 --------------- = 21*19*17*3 5 *4*3*2*1

OpenStudy (anonymous):

5*(21*20*19*18*17) = 12209400